Sara wants to buy a house and needs $35,560 for a down payment. She currently has $28,000 in a savings account that earns 9% simple interest. How long must Sara keep the money in the savings account in order to have enough for the down payment on the house?
Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

P=Po(1+.09)^t
35,560/28,000=1.09^t=1.27
ln of both sides
ln (1.27)=t*ln (1.09)
t=ln(1.27)/ln(1.09)=2.77 years

Answer by MathTherapy(10552)

Sara wants to buy a house and needs $35,560 for a down payment. She currently has $28,000 in a savings account that earns 9% simple interest. How long must Sara keep the money in the savings account in order to have enough for the down payment on the house?
SI = PTR
7,560 = 28,000 * T * .09
7,560 = 2,520T
T, or time =
If the interest that’s earned each year is WITHDRAWN EVERY YEAR, right after being added to the account, then it’d be 3 years, since
at all times, only the $28,000 would be earning interest. Otherwise, it'd be COMPOUNDED, and time would be shorter, at 2.77 years.
Who formulates these problems?